OpenStudy (anonymous):

Evaluate the function f(x) = 3x − 2, when x = −1

OpenStudy (trojanpoem):

replace each x with -1 note f(x) is equivalent to saying y f(-1) = (3* -1) -2 = -5

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x) = 3x - 2 \\[0.9em] }\) if I wanted to find \(\large\color{black}{ \displaystyle f(2) }\) I would do the following \(\large\color{black}{ \displaystyle :\\[1.0em] }\) \(\large\color{black}{ \displaystyle f(x) = 3x - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{2}) = 3(\color{blue}{2}) - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{2}) = 6 - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{2}) = 4 \\[0.3em] }\) if I wanted to find \(\large\color{black}{ \displaystyle f(-5) }\) I would do the following \(\large\color{black}{ \displaystyle :\\[1.0em] }\) \(\large\color{black}{ \displaystyle f(x) = 3x - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{-5}) = 3(\color{blue}{-5}) - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{-5}) = -15 - 2 \\[0.3em] }\) \(\large\color{black}{ \displaystyle f(\color{blue}{-5}) = -17 \\[0.3em] }\)

OpenStudy (solomonzelman):

these are just examples. YOu needed to do the same exact thing, but for \(\large\color{black}{ \displaystyle f(-1) }\).